OpenStudy (anonymous):

What is the equation for a line that goes through points (-8,5)and(4,-7)?

OpenStudy (zale101):

m=(-7-5)/(4-(-8))= -1 equation: y-5=(-1)[x-(-8)] y-5=(-1)[x+8] Solve it and try to get the equation...